OpenStudy (nthenic_oftime):

you will get a medal :) Locate the foci of the ellipse. Show your work. x^2/36+y^2/11=1

OpenStudy (anonymous):

any ideas where the ellipse center is at?

OpenStudy (anonymous):

notice your center \(\cfrac{(x-{\color{brown}{ h}})^2}{{\color{purple}{ a}}^2}+\cfrac{(y-{\color{blue}{ k}})^2}{{\color{purple}{ b}}^2}=1 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad vertices\ ({\color{brown}{ h}}\pm a, {\color{blue}{ k}}) \\ \quad \\ % ellipse, vertical major axis \cfrac{(x-{\color{brown}{ h}})^2}{{\color{purple}{ b}}^2}+\cfrac{(y-{\color{blue}{ k}})^2}{{\color{purple}{ a}}^2}=1 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad vertices\ ({\color{brown}{ h}}, {\color{blue}{ k}}\pm a)\)

OpenStudy (anonymous):

anyhow, get the center, see if it's horizontal or a vertical major axis then get the "c" distance, "c" is the distance from the center to either foci and \(\bf c=\sqrt{a^2-b^2}\)

Directrix (directrix):

Continuing from above, c = sqrt(36 - 11) c = sqrt(25) c = 5 @nthenic_oftime Can you get the foci coordinates now.